OpenStudy (babyslapmafro):

Please help me simplify the following. 10th root of 32a^5

OpenStudy (babyslapmafro):

\[\sqrt[10]{32a ^{5}}\]

ganeshie8 (ganeshie8):

hint : 32 = 2^5

OpenStudy (anonymous):

\[\sqrt[10]{32a^5}=\sqrt[10]{2*2*2*2*2*a^5}=\sqrt[10]{2^5a^5}=\sqrt[10]{(2a)^5}=((2a)^5)^{1/10}=(2a)^{5/10}=(2a)^{1/2}=\sqrt{2a}\]

OpenStudy (ksaimouli):